1) The combustion of 3.42g of a compound is known go contain only nitrogen and hydrogen gave 9.82g of NO2 and 1.85g of water. Determine the empirical formula of this compound and show work.

2) A compound with a known molecular mass (146.99g/mol.) that contain only C,H and Cl was studies by combustion analysis. When a 0.367 g was combusted. 0.0659g of CO2 and 0.0892g of water formed. What are the empirical and molecular formulas of the compound?
1
Expert's answer
2017-10-09T14:46:06-0400
1) The equation of combustion:
NxHy+(x+y/4)O2 → xNO2+y/2H2O
m(N) = 14/(14+16*2) * 9,82 = 2,989 (grams)
m(H) = 2/(2+16)*1,85 = 0,206 (grams)
m(N)/m(H) = 2,989/0,206=14,51
14x/1y=14,51
x/y=1,036≈1
Empirical formula of unknown compound: N1H1

2) 2) Empirical formula: C3H2Cl

Molecular formula: C6H4Cl2